OpenStudy (anonymous):

int 2x dx from 2 to 3

OpenStudy (anonymous):

\[\int\limits_{2}^{3} 2x dx\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Thanks!

OpenStudy (unklerhaukus):

\[\int\limits_{2}^{3} 2x \cdot\text dx\]\[=2\int\limits_{2}^{3} x \cdot\text dx\]\[=2\left. \frac {x^2}{2} \right|_{2}^{3}\]\[=2 \frac {3^2-2^2}{2} \]